OpenStudy (anonymous):

please solve: Linear Equations by substitution. (with solution) x=3y+4 2x-8y-6=0

OpenStudy (anonymous):

2(3y+4)-8y-6 = 0 6y + 8 -8y -6 =0 -2y + 2 = 0 2 = 2y y = 1 x =3y +4 x = 3(1) + 4 X = 3 + 4 x =7